3. Write a system of two equations in two unknowns. Solve the system by using the substitution method. « The length of a rectangular field is 12 feet longer than the width. If the perimeter is 144 feet, then what are the length and width ? » Show your work.
Answer by ankor@dixie-net.com(22740)

Write a system of two equations in two unknowns. Solve the system by using the substitution method. «
:
"The length of a rectangular field is 12 feet longer than the width." says:
L = W + 12
:
"If the perimeter is 144 feet," says
2L + 2W = 144
:
then what are the length and width ? » Show your work.
:
Using the 1st equation, substitute (W+12) for L in the 2nd equation:
2(W+12) + 2W = 144
:
Simplify, divide equation by 2 and you have:
W + 12 + W = 72
:
2W = 72 - 12
:
W = 60/2
:
W = 30
:
I'll let you find the length, check both solutions in the perimeter equation